OpenStudy (anonymous):

Write an equation for the tangent line to f(x) = -4x + 3/x at x = -1 How do i solve this?

OpenStudy (anonymous):

So, we need to know the slope at this point, this is given by: \[ \frac{d}{dx}\left(-4x+\frac{3}{x}\right)=-4-\frac{3}{x^2} \]Evaluating this at \(x=-1\): \[ -4-\frac{3}{x^2}\Bigg|_{x=-1}=-4+3=-1 \]So the slope of the line is -1 at this point. Now, find a point on the line: \[ -4(-1)+\frac{3}{-1}=-7=y \]Which means: \[ y+7=-1(x+1) \]Et voilá.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

Which would be the correct answer: A) -7x + y = -6 B) -6x + 2y = -5 C) 7x + y = -6 D) 6x + 2y = -5

OpenStudy (anonymous):

Wait, I screwed up, it's: \[ -4-\frac{3}{x^2}\Bigg|_{x=-1}=-4-3=-7 \]So: \[ y+7=-7(x+1) \]

OpenStudy (anonymous):

oh! that explains why i coudnt get the answer. Thanks for the assistance

OpenStudy (anonymous):

Sure thing

OpenStudy (anonymous):

One more question..How would i get the end results in one of those 4 possible results i mentioned? can you show me?

OpenStudy (anonymous):

Just solve for the same form, it's roughly the same case.